john got 3 books and 5 pens for 60 dollars each book cost 5 times as much as a pen,find the cost of the two books and seven pens
Answer by checkley79(3341)

3B+5P=60
B=5P
3*5P+5P=60
15P+5P=60
20P=60
P=60/20
P=$3.00 PRICE OF EACH PEN.
B=5*3=$15.00 PRICE OF EACH BOOK.
2B+7P=?
2*15+7*3=30+21=$51.00 COST OF 2 BOOKS & 7 PENS.